# From User Preferences to Base Score Extraction Functions in Gradual Argumentation

### v1

@@ Abstract.
Gradual argumentation is a field of symbolic AI which is attracting attention for its ability to support transparent and contestable AI systems. It is considered a useful tool in domains such as decision-making, recommendation, debate analysis, and others. The outcomes in such domains are usually dependent on the arguments’ base scores, which must be selected carefully. Often, this selection process requires user expertise and may not always be straightforward. On the other hand, organising the arguments by preference could simplify the task. In this work, we introduce Base Score Extraction Functions, which provide a mapping from users’ preferences over arguments to base scores. These functions can be applied to the arguments of a Bipolar Argumentation Framework (BAF), supplemented with preferences, to obtain a Quantitative Bipolar Argumentation Framework (QBAF), allowing the use of well-established computational tools in gradual argumentation. We outline the desirable properties of base score extraction functions, discuss some design choices, and provide an algorithm for base score extraction. Our method incorporates an approximation of non-linearities in human preferences to allow for better approximation of the real ones. Finally, we evaluate our approach both theoretically and experimentally in a robotics setting, and offer recommendations for selecting appropriate gradual semantics in practice.
@@ 1. Introduction
Abstract Argumentation Frameworks (AFs) Dung_JLP95 have long been known to be a promising tool for adaptable decision-making systems Amgoud_AAI09. These formalisms are composed of arguments and relations between them, representing conflicts. Semantics, whether extension-based or gradual, may then be used to determine the acceptance of the arguments, e.g., as potential decisions to be taken. However, in real-world scenarios involving human reasoning, having only arguments and their relations is rarely sufficient to align the decision towards human intentions. For example, some studies claim that the acceptability of arguments is related to the preferences of the audience to which they are addressed Kaci_CMA18.
Aligning decision-making systems with human preferences Gabriel_MM20, and also values Sierra_aamas19, has never been more important, given their increasing prevalence in humans’ daily lives. Within the field of Computational Argumentation, this has been accomplished by extending AFs into Preference-based AFs (PAFs) Amgoud_uai13, in which the arguments’ acceptability is based not only on their relations and the chosen semantics, but also on a preference ordering.
Figure 1. We introduce a methodology for converting a BAF, supplemented with a preference ordering of the arguments (top), into a QBAF (bottom). From the set of arguments and a user’s preference ordering, a Base Score Extraction Function $\nu$ is applied to the preference ordering to obtain the arguments’ base scores $\tau$ (middle), which allow for the QBAF to be used as normal for decision-making, i.e. comparing the decision arguments and selecting that with the highest strength. The output of the decision-making system is personalised to the user’s preferences.
@@
A current limitation in extension-based PAFs is that preferences are used to reduce or modify the framework Amgoud_IJAR14; Kaci_CMA18. Those reductions correspond to different intuitions, and each one is subject to criticism. In Battaglia_icsum24, reductions are avoided by extending PAFs to Quantitative Bipolar AFs (QBAFs) Baroni_IJAR19. In QBAFs, the arguments can attack and support each other, and each argument is assigned a base score, i.e. an intrinsic strength. Gradual semantics of QBAFs then assign a final strength to each argument, providing a different form of information than traditional AFs. QBAFs are currently used in several domains, such as decision support Battaglia_icsum24, product recommendation Rago_ijcai18, review aggregation Cocarascu_aamas19; Rago_AI25, decision-making reasoning Freedman_aaai25, given their potential for giving transparent and contestable systems Leofante_KR24; Yin_arxiv25. Nonetheless, in some scenarios, it is often unclear how to set the arguments’ base scores.
In this work, we propose a theoretically and experimentally motivated methodology for generating these base scores based on a given set of user preferences. Our approach is based on adding preferences to Bipolar Argumentation Frameworks (BAFs) Cayrol_ecsqaru05, i.e., AFs with an additional relation of support (but without a base score). We operate on the hypothesis that the given preference ordering allows for a base score extraction function to predict the arguments’ base scores. Although the focus of the paper is on converting BAFs into QBAFs, it could be applicable to AFs and support argumentation frameworks. Our approach is represented in Fig. 1. To ensure that our approach is able to represent realistic user preferences, on top of the usual I prefer $a$ over $b$, we introduce the possibility of having a preference relation in which one argument is much more preferred than another. We then provide some general axioms that any base score extraction function should satisfy. Next, we provide different design choices for adapting the base score extraction, which are formally characterised as potentially desirable, optional properties. Finally, we introduce some concrete base score extraction functions and evaluate them through both a theoretical analysis and experimentation on a proposed use case, assistive feeding in robotics.
Our contributions are as follows: (i) the introduction of base score extraction functions, for extracting the base scores of a set of arguments given a preference ordering, extended to non-linear preferences; (ii) the definition of a set of axioms and properties for base score extraction functions; (iii) the introduction of two concrete base score extraction functions; (iv) a theoretical and experimental evaluation of the two base score extraction functions.
@@ 2. Preliminaries
As mentioned, this work uses QBAFs for decision-making. In this context, the possible options of the decision process are included in the QBAF. These are called decision arguments. The QBAF is represented as a set of trees, with the root of each tree corresponding to a decision argument Rago_kr23. Let $\mathcal{Q}$ be a QBAF $\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+},\tau\rangle$. For any arguments $a,b\in\mathcal{X}$, let a path from $a$ to $b$ be defined as a sequence of relations $(\gamma_{0},\gamma_{1}),$..., $(\gamma_{n-1},\gamma_{n})$ of length $n>0$, where $\gamma_{0}=a$ and $\gamma_{n}=b$, and, for any $1\leq i\leq n,(\gamma_{i-1},\gamma_{i})\in\mathcal{R}^{+}\cup\mathcal{R}^{-}$. Then, given a set of decision arguments $D\subseteq\mathcal{X}$, $\mathcal{Q}$ is a QBAF for $D$ iff i) $\nexists a\in\mathcal{X}\setminus\{D\}$ such that $\exists d\in D$ where $(d,a)\in\mathcal{R}^{+}\cup\mathcal{R}^{-}$ ii) $\forall a\in\mathcal{X}\setminus\{D\}$ there is a path from $a$ to at least one $d\in D$; and iii) $\nexists a\in\mathcal{X}$ with a path from $a$ to $a$. We leave cyclic frameworks to future work, as in Potyka_aaai21; Rago_kr23; Rago_AI25.
As a running example, we will use robot-assisted feeding. Here, a robot feeds a user who is unable to eat independently and must decide the pace of feeding, either slow or fast. The robot uses a QBAF drawing arguments from both the user and the robot’s observations, e.g., user status and history. The decision is made by selecting the decision argument with the higher strength. The arguments are: $(a)$ The patient gets bored if eating slowly, $(b)$ eating slowly does not stress the patient out, $(c)$ the patient is vulnerable and stress must be avoided, $(d)$ if the user does not eat fast, they will not have time with their niece who is coming to visit, $(e)$ the patient wants to tell their niece something important, and doing so will help the patient to feel more relaxed, and $(f)$ if the patient eats quickly, there is a (low) risk of dysphagia. The decision arguments are slow ($D_{1}$) and fast ($D_{2}$), which are possible feeding paces the robot can perform. The option selected is the one with the highest final strength. The framework is represented in Fig. 1.
@@ 3. Extracting Base Scores from Preference Orderings
In this section, we show that adding preferences to BAFs allows us to define a base score extraction function, which may exhibit some defined desirable axioms and properties. Then, the base score extraction function is used to extract the arguments’ base scores, effectively converting the BAF into a QBAF.
@@ 3.1. Adding Preferences to BAFs
Adding preferences over arguments is a common approach to personalising argumentation frameworks for users Modgil_AI13; Amgoud_uai13; Mailly_kr20. We introduce the formal definition of preferences over arguments:
@@
Given a set of arguments $\mathcal{X}$, a preference $\succeq_{p}$ over $X$ is a reflexive and transitive relation on $\mathcal{X}$. Given $\in\mathcal{X}, aba\succeq b$ and $a\npreceq b$ will be denoted as $a\succ b$, while $a\preceq b$ and $a\succeq b$ will be denoted as $a\simeq b$.
Setting the base scores according to the preference ordering of a BAF’s arguments in a decision-making environment can give an intuition of the most suitable decision in line with a user’s preferences. The following examples show the importance of properly setting the arguments’ base scores:
@@
Consider the framework from Fig. 1. If the base scores are arbitrarily set to 0.5, giving equal importance to all arguments, using the Quadratic Energy Potyka_kr18 model, the final strengths of the decision arguments are $\sigma(D_{1})=0.5$ for slow, and $\sigma(D_{2})=0.505$ for fast, then the robot would choose to move fast.
Now imagine a scenario where a user gives more importance to the arguments related to its safety ($c$ and $f$), less importance to the arguments related to its relaxation ($b$ and $e$), and the least importance to the arguments related to its enjoyment ($a$ and $d$):
@@
(cont) The resulting preference ordering is: $c\simeq f\succ b\simeq e\succ a\simeq d$. Intuitively, the base scores of $c$ and $f$ should be greater than those of $b$ and $e$, and these greater than the base scores of $a$ and $d$. For example, they could be set to $\tau(c)=\tau(f)=0.75$, $\tau(b)=\tau(e)=0.5$, and $\tau(a)=\tau(d)=0.25$. Here, using the Quadratic energy model, the final strengths are $\sigma(D_{1})=0.54$ and $\sigma(D_{2})=0.45$, then the robot would choose to move slowly, which is the opposite of what has been selected with equal base scores.
@@ 3.2. Adapting to Non-Linear Human Preferences
Human preference relations are generally not linear Northcraft_OBHDP98; people’s choices and priorities often fluctuate instead of increasing or decreasing in a simple, linear manner as conditions change. This non-linearity makes it complex to set the base scores properly. A possible approach to including these non-linearities could be achieved by allowing the user to order their preferences gradually within a range. However, setting a gradual score is usually slower and has more sample noise than cardinal ordering Yannakakis_acii11; Shah_arxiv14.
Another approach, which is used in this work, is to approximate the non-linearities by introducing a new preference relation, where a user may have a much greater ($\succ\!\succ$) preference for one argument over another. This novel relation considers the non-linearity and is easy to recognise from human feedback. Definition 1 is extended as follows:
@@
Given $a,b\in\mathcal{X}$, if $a$ is much more preferred than $b$, the preference relation will be denoted as $a\succ\!\succ b$. This relation is transitive and satisfies $a\succ b$. Therefore, a preference relation between arguments $a,b\in\mathcal{X}$ in a preference ordering $\succeq_{p}$ can be:
@@
Given the arguments $a,b,c\in\mathcal{X}$, with a preference ordering: $a\succ\!\succ b\succ c$. Here, we may expect that $\tau(a)$ is much greater than $\tau(b)$ and $\tau(c)$, e.g., 0.9, 0.3, and 0.1, respectively.
@@ 3.3. Base Score Extraction Functions
Therefore, we introduce base score extraction functions that compute the arguments’ base scores given a preference ordering.
@@
A Base Score Extraction Function $\nu$ is a function that takes a set of arguments $\mathcal{X}$ and a set of preference orderings $\mathcal{P}$ to obtain a set of all the base score functions $\mathcal{T}$, therefore $\nu:\mathcal{X}\times\mathcal{P}\rightarrow\mathcal{T}$.
Obtaining the base scores of a set of arguments allows the transformation of BAFs into QBAFs. To simplify the formalisations in the following sections, we consider $\nu_{P}(\mathcal{X})$ as the base score extraction function $\nu$ applied to $\mathcal{X}$ with a given preference ordering $\succeq_{p}$.
@@ 3.4. Desirable Properties of Base Score Extraction Functions
In the following, we present the different axioms and properties of the base score extraction functions. First, we define the Preference Coherence axiom, which states that if one argument is preferred over another, the former’s base score will be greater.
@@
(Preference Coherence) Given a set of arguments $\mathcal{X}$ with arguments $a,b\in\mathcal{X}$ and a preference ordering $\succeq_{p}$, a base score extraction function $\nu_{p}(\mathcal{X})$ satisfies Preference Coherence iff:
@@
(Preference Relation Coherence) Given a set of arguments arguments $a,b,c,d\in\mathcal{X}$ and the preference ordering $a\succ\!\succ b\succ c\simeq d$, a base score extraction function $\nu$ satisfies Preference Relation Coherence iff $\tau(a)-\tau(b)>\tau(b)-\tau(c)>\tau(c)-\tau(d)$.
@@
Let $\succeq_{p}$ and $\succeq_{p^{\prime}}$ be two preference orderings over two identical sets of arguments, $\mathcal{X}$ and $\mathcal{X}^{\prime}$, respectively. Those preference orderings are isomorphic iff there is a bijective function $f:\mathcal{X}\to\mathcal{X^{\prime}}$ such that $\forall i,j\in\mathcal{X},\ i\succeq_{p}j\Leftrightarrow f(i)\succeq_{p^{\prime}}f(j)$ [^1] We omit the subscript of the preference relation when unambiguous.. It is denoted as $\succeq_{p}\simeq\succeq_{p^{\prime}}$.
We require that a base score extraction function preserves the structural distinctions of users’ preferences. If two users have different (non-isomorphic) preference orderings, the assigned base scores must reflect these differences.
@@
(Preference Structure Coherence) Given two preference orderings that are not isomorphic $\succeq_{p}\not\simeq\succeq_{p^{\prime}}$, a base score extraction function $\nu$ satisfies Preference Structure Coherence iff for a set of arguments $\mathcal{X}$, then $\nu_{p}(\mathcal{X})\neq\nu_{p^{\prime}}(\mathcal{X})$.

### v3

@@ Abstract.
Gradual argumentation is a sub-field of Computational Argumentation from symbolic AI which is attracting attention for its ability to support transparent and contestable AI systems. It is considered a useful tool in domains such as decision-making, recommendation, debate analysis, amongst others. The outcomes in such domains are usually dependent on the arguments’ base scores, which must be selected carefully. Often, this selection process requires user expertise and may not always be straightforward. On the other hand, organising the arguments by preference could simplify the task. In this work, we introduce Base Score Extraction Functions, which provide a mapping from users’ preferences over arguments to base scores. These functions can be applied to the arguments of a Bipolar Argumentation Framework (BAF), supplemented with preferences, to obtain a Quantitative Bipolar Argumentation Framework (QBAF), allowing the use of well-established computational tools in gradual argumentation. We outline the desirable properties of Base Score Extraction Functions, discuss some design choices, and provide an algorithm for base score extraction. Our method incorporates an approximation of non-linearities in human preferences to allow for better approximation of the real ones. Finally, we evaluate our approach both theoretically and experimentally in a robotics setting, and offer recommendations for selecting appropriate gradual semantics in practice.
@@ 1. Introduction
Abstract Argumentation Frameworks (AFs) Dung95 have long been known to be a promising method for adaptable decision-making systems Amgoud_AAI09. These frameworks are composed of arguments and relations between them, representing conflicts. Semantics, whether extension-based Dung95 or gradual AAgradual19, may then be used to determine the acceptance of the arguments, e.g., as potential decisions to be taken. However, in real-world scenarios involving human reasoning, having only arguments and their relations is rarely sufficient to align the decision towards human intentions. For example, some studies claim that the acceptability of arguments is related to the preferences of the audience to which they are addressed Kaci_CMA18.
Aligning decision-making systems with human preferences Gabriel_MM20, and also values Sierra_aamas19, has never been more important, given their increasing prevalence in humans’ daily lives. Within the field of Computational Argumentation, this has been accomplished by extending AFs into Preference-based AFs (PAFs) Amgoud_uai13, in which the arguments’ acceptability is based not only on relations and the chosen semantics, but also on a preference ordering.
Figure 1. We introduce a methodology for converting a BAF supplemented with a preference ordering of the arguments (top), into a QBAF (bottom). A Base Score Extraction Function $\nu$ is applied to the preference ordering to obtain the arguments’ base scores $\tau$ (middle), which allow for the QBAF to be used as normal for decision-making. The output of the decision-making system is adapted to the user’s preferences.
@@
A current limitation in extension-based PAFs is that preferences are used to reduce or modify the framework Amgoud_IJAR14; Kaci_CMA18. Those reductions correspond to different intuitions, and each one is subject to criticism. Moreover, many decision-making settings require a single output, whereas qualitative frameworks may return multiple undominated options or none at all Amgoud_AI09. In contrast, Gradual Argumentation ensures a determinate outcome while remaining transparent about score contributions Civit_arxiv25. In Battaglia_icsum24, reductions are avoided by extending PAFs to Quantitative Bipolar AFs (QBAFs) Baroni_IJAR19. In QBAFs, the arguments can attack and support each other, and each argument is assigned a base score, i.e. an intrinsic strength. Gradual semantics of QBAFs Baroni_IJAR19 then assign a final strength to each argument, providing a different form of information than traditional AFs. QBAFs are currently used in several domains, such as decision support Battaglia_icsum24, product recommendation Rago_ijcai18, review aggregation Cocarascu_aamas19; Rago_AI25, and decision-making Freedman_aaai25, given their potential for giving transparent and contestable systems Leofante_KR24; Yin_arxiv25. Nonetheless, in some scenarios, it is often unclear how to set the arguments’ base scores.
In this work, we propose a theoretically and experimentally motivated methodology for generating these base scores based on a given set of user preferences. Our approach is based on adding preferences to Bipolar Argumentation Frameworks (BAFs) Cayrol_ecsqaru05, i.e., AFs with an additional relation of support (but without a base score). We operate on the hypothesis that the given preference ordering allows for a Base Score Extraction Function (BSEF) to predict the arguments’ base scores. Our approach is illustrated in Fig. 1. Note that, to ensure that it is able to represent realistic user preferences, on top of the usual ‘I prefer one argument over another’, we introduce the possibility of having a preference relation in which one argument is much more preferred than another. We then give some general axioms that any BSEF should satisfy. Next, we provide different design choices for adapting the base score extraction, which are formally characterised as potentially desirable, optional properties. Finally, we introduce some concrete BSEFs and evaluate them through both a theoretical analysis and experimentation on a proposed use case, namely assistive feeding in robotics.
Our contributions are as follows: (i) the introduction of BSEFs, which extract the base scores of a set of arguments given a preference ordering, extended to non-linear preferences; (ii) the definition of a set of axioms and properties for BSEFs; (iii) the introduction of two concrete BSEFs; and (iv) a theoretical and experimental evaluation of the two BSEFs.
@@ 2. Preliminaries
As mentioned, we use QBAFs for decision-making. In this context, the possible options of the decision process are included in the QBAF. These are called decision arguments. The QBAF is represented as a set of trees, with the root of each tree corresponding to a decision argument Rago_kr23. (We leave cyclic QBAFs to future work, as in Potyka_aaai21; Rago_kr23; Rago_AI25.) Let $\mathcal{Q}$ be a QBAF $\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+},\tau\rangle$. For any arguments $a,b\in\mathcal{X}$, let a path from $a$ to $b$ be defined as a sequence $(\gamma_{0},\gamma_{1}),$..., $(\gamma_{n-1},\gamma_{n})$ of length $n>0$, where $\gamma_{0}=a$ and $\gamma_{n}=b$, and, for any $1\leq i\leq n,(\gamma_{i-1},\gamma_{i})\in\mathcal{R}^{+}\cup\mathcal{R}^{-}$. Then, given a set of decision arguments $D\subseteq\mathcal{X}$, $\mathcal{Q}$ is a QBAF for $D$ iff i) $\nexists a\in\mathcal{X}\setminus\{D\}$ such that $\exists d\in D$ where $(d,a)\in\mathcal{R}^{+}\cup\mathcal{R}^{-}$ ii) $\forall a\in\mathcal{X}\setminus\{D\}$ there is a path from $a$ to at least one $d\in D$; and iii) $\nexists a\!\in\!\mathcal{X}$ with a path from $a$ to $a$. We set the base score of the decision arguments to 0.5.
In the remainder of the paper, unless otherwise specified, we will assume a BAF $\langle\mathcal{X},\mathcal{R}^{-},\mathcal{R}^{+}\rangle$.
As a running example, we will use robot-assisted feeding. Here, a robot feeds a user who cannot eat independently and must decide the pace of feeding, either slow or fast. The robot uses a BAF drawing arguments from both the user and the robot’s observations, e.g., user status and history. The arguments are: $(a)$ Eating slowly causes boredom; $(b)$ eating slowly does not stress the patient out; $(c)$ the patient is vulnerable and stress must be avoided; $(d)$ eating slowly reduces time with a visiting niece; $(e)$ the patient wants to tell their niece something important, and doing so will help the patient to feel more relaxed; and $(f)$ eating quickly carries a (low) risk of dysphagia. The decision arguments are slow ($D_{1}$) and fast ($D_{2}$). The BAF is represented in Fig. 1. The aim is to select the best option, namely any decision argument with the highest final strength in the QBAF obtained after predicting the arguments’ base scores to adhere to the preferences (again, as in Fig. 1).
@@ 3. Extracting Base Scores from Preference Orderings
In this section, we show that adding preferences to BAFs allows us to define a BSEF, which may exhibit some defined desirable axioms and properties. Then, the BSEF is used to extract the arguments’ base scores, effectively converting the BAF into a QBAF.
@@ 3.1. Adding Preferences to BAFs
Adding preferences over arguments is a common approach to personalising argumentation frameworks for users Modgil_AI13; Amgoud_uai13; Mailly_kr20. We formally define preferences over arguments as follows:
@@
A preference $\succeq$ over $\mathcal{X}$ is a reflexive and transitive relation on $\mathcal{X}$. Given $\in\mathcal{X}, aba\succeq b$ and $a\npreceq b$ will be denoted as $a\succ b$, while $a\succeq b$ and $a\preceq b$ will be denoted as $a\simeq b$.
Setting the base scores according to the preference ordering of a BAF’s arguments in a decision-making environment can give an intuition of the most suitable decisions in line with a user’s preferences. The following examples show the importance of properly setting the arguments’ base scores:
@@
Consider the framework from Fig. 1. If the base scores are arbitrarily set to 0.5, giving equal importance to all arguments, using the Quadratic Energy (QE) Model Potyka_kr18, the final strengths of the decision arguments are $\sigma(D_{1})=0.5$ for slow, and $\sigma(D_{2})=0.505$ for fast. Then the robot would choose to move fast.
@@
(Example 3.1 Cont.) Imagine a scenario where a user gives more importance to the arguments related to its safety (e.g., $c$ and $f$), less importance to those related to its relaxation (e.g., $b$ and $e$), and the least importance to the arguments related to its enjoyment (e.g., $a$ and $d$). The resulting preference ordering is: $c\simeq f\succ b\simeq e\succ a\simeq d$. Intuitively, the base scores of $c$ and $f$ should be greater than those of $b$ and $e$, and which should in turn be greater than the base scores of $a$ and $d$. For example, they could be set to $\tau(c)=\tau(f)=0.75$, $\tau(b)=\tau(e)=0.5$, and $\tau(a)=\tau(d)=0.25$. Here, using the QE Model again, the final strengths are $\sigma(D_{1})=0.54$ and $\sigma(D_{2})=0.45$. Then the robot would choose to move slowly, which is the opposite of what was selected earlier.
@@ 3.2. Adapting to Non-Linear Human Preferences
Human preference relations are generally not linear Northcraft_OBHDP98. People’s choices and priorities often fluctuate instead of increasing or decreasing in a simple, linear manner as conditions change. This non-linearity makes it complex to set the base scores properly. A possible approach to including these non-linearities could be achieved by allowing the user to order their preferences gradually within a range. However, setting a gradual score is usually slower and has more sample noise than cardinal orderings Yannakakis_acii11; Shah_arxiv14.
In this work, we approximate the non-linearities by introducing a new preference relation, where a user may have a much greater ($\succ\!\succ$) preference for one argument over another. This relation considers the non-linearity and is easy to recognise from human feedback. Definition 1 is extended as follows:
@@
Given $a,b\in\mathcal{X}$, if $a$ is much more preferred than $b$, the preference relation will be denoted as $a\succ\!\succ b$. This relation is transitive, and $a\succ\!\succ b$ implies $a\succ b$. Therefore, a preference relation between arguments $a,b\in\mathcal{X}$ in a preference ordering $\succeq$ can be:
  • $a\simeq b$, which denotes indifference;
@@
Given the arguments $a,b,c\in\mathcal{X}$, with a preference ordering: $a\succ\!\succ b\succ c$, we may expect that $\tau(a)$ is much greater than $\tau(b)$ and $\tau(c)$, e.g., 0.9, 0.3, and 0.1, respectively.
@@ 3.3. Base Score Extraction Functions
Therefore, we introduce BSEFs that compute the arguments’ base scores given a preference ordering.
@@
A BSEF $\nu$ is a function that maps a set of preference orderings $\mathcal{P}$ to a set of base score functions $\mathcal{T}$ (for arguments in $\mathcal{X}$) ; namely $\nu:\mathcal{P}\rightarrow\mathcal{T}$.
Obtaining the base scores of a set of arguments allows the transformation of BAFs into QBAFs.
@@ 3.4. Desirable Properties of Base Score Extraction Functions
In the following, we present the different axioms and properties of the BSEFs. First, we define the Preference Coherence axiom, which states that if one argument is preferred over another, the former’s base score will be greater.
@@
(Preference Coherence) Given a preference ordering $\succeq$, a BSEF $\nu(\succeq)=\tau$ satisfies Preference Coherence iff, for any arguments $a,b\in\mathcal{X}$:
@@
(Preference Relation Coherence) Given a set of arguments arguments $a,b,c,d\in\mathcal{X}$ and the preference ordering $a\succ\!\succ b\succ c\simeq d$, a BSEF $\nu(\succeq)=\tau$ satisfies Preference Relation Coherence iff $\tau(a)-\tau(b)>\tau(b)-\tau(c)>\tau(c)-\tau(d)$.
@@
Let $\mathcal{X}^{\prime}$ be a set of arguments and $\succeq$ and $\succeq^{\prime}$ be two preference orderings over $\mathcal{X}$ and $\mathcal{X}^{\prime}$, respectively. Those orderings are isomorphic, denoted $\succeq\simeq\succeq^{\prime}$, iff there is a bijective function $f:\mathcal{X}\to\mathcal{X}^{\prime}$ such that $\forall i,j\in\mathcal{X},\ i\succeq j\Leftrightarrow f(i)\succeq^{\prime}f(j)$.
We require that a BSEF preserves the structural distinction of users’ preferences. If two users have different (non-isomorphic) preference orderings, the assigned base scores must reflect it.
@@
(Preference Structure Coherence) Given two preference orderings $\succeq$ and $\succeq^{\prime}$ (over the same $\mathcal{X}$) that are not isomorphic ($\succeq\not\simeq\succeq^{\prime}$), then a BSEF $\nu$ satisfies Preference Structure Coherence iff $\nu(\succeq)\!\neq\!\nu(\succeq^{\prime})$.